### v1

# Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning

Abstract
Teaching large language models (LLMs) to be faithful in the provided context is crucial for building reliable information-seeking systems. Therefore, we propose a systematic framework, Canoe, to improve the faithfulness of LLMs in both short-form and long-form generation tasks without human annotations. Specifically, we first synthesize short-form question-answering (QA) data with four diverse tasks to construct high-quality and easily verifiable training data without human annotation. Also, we propose Dual-GRPO, a rule-based reinforcement learning method that includes three tailored rule-based rewards derived from synthesized short-form QA data, while simultaneously optimizing both short-form and long-form response generation. Notably, Dual-GRPO eliminates the need to manually label preference data to train reward models and avoids over-optimizing short-form generation when relying only on the synthesized short-form QA data. Experimental results show that Canoe greatly improves the faithfulness of LLMs across 11 different downstream tasks, even outperforming the most advanced LLMs, e.g., GPT-4o and OpenAI o1.111 The data, code, and models will be available at https://github.com/S1s-Z/CANOE.
[…]
tcb@breakable
Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning
Shuzheng Sithanks:  Equal Contribution.♠♢, Haozhe Zhao11footnotemark: 1, Cheng Gao11footnotemark: 1, Yuzhuo Bai, Zhitong Wang Bofei Gao, Kangyang Luo, Wenhao Li, Yufei Huang, Gang Chen Fanchao Qi, Minjia Zhang, Baobao Chang, and Maosong Sun Tsinghua University   Peking University   DeepLang AI University of Illinois Urbana-Champaign
1 Introduction
Recent progress in large language models (LLMs) has revolutionized text generation with their remarkable capabilities (OpenAI, 2023; DeepSeek-AI et al., 2025b). LLMs are widely used to generate fluent and coherent text responses based on the provided contextual information, e.g., document question answering (QA) (Wang et al., 2024) and text summarization (Zhang et al., 2024). However, LLMs often generate responses that are not faithful or grounded in the input context, i.e., faithfulness hallucinations (Ji et al., 2023; Huang et al., 2024; Si et al., 2025), which can undermine their trustworthiness. Maintaining faithfulness to the context is especially important in fields where accurate information transfer is essential (Duong et al., 2025). For instance, in legal summarization (Dong et al., 2025), the text output must reflect the content of legal documents without introducing any distortions.
Figure 1: Average score on 11 downstream tasks vs model size. With only 7B parameters, Canoe already exceeds state-of-the-art LLMs like GPT-4o and o1.
However, improving the faithfulness of LLMs faces three key challenges. Specifically, (1) Faithfulness is difficult to improve by simply scaling model parameters: Previous works (Xie et al., 2024; Li et al., 2025) find that LLMs may overly rely on internal knowledge learned from extensive pre-training data while disregarding provided contexts, i.e., the knowledge conflicts (Xu et al., 2024b). When the model parameters increase and internal knowledge grows, this may lead to greater knowledge conflicts and further lower the faithfulness of LLMs (Ming et al., 2025). Thus, it is necessary to explore the tailored post-training method to improve the faithfulness instead of simply scaling the model parameters. (2) Faithfulness is challenging to consistently boost across different downstream tasks: Recently, several methods (Li et al., 2024; Duong et al., 2025) have been proposed to improve the faithfulness of LLMs for different tasks. For example, Bi et al. (2024) aligns LLMs through DPO (Rafailov et al., 2023) with constructed faithful and unfaithful short-form completions, improving the performance of LLMs on short-form QA tasks. However, these recent methods are designed for specific tasks, so they fail to consistently improve the faithfulness of LLMs across various tasks, like text summarization and multiple-choice questions, because these tasks can vary greatly. (3) Data used to enhance faithfulness is hard to scale: This issue is especially problematic with data used to improve the faithfulness in long-form generation tasks. Unlike tasks with clear answers, e.g., short-form fact-seeking QA tasks (Wei et al., 2024), there is no standard way to ensure data quality in long-form generation tasks (Duong et al., 2025). Thus, data is typically annotated by humans (Kryscinski et al., 2020; Zhu et al., 2023), which is costly and not scalable.
To tackle these challenges, we propose a systematic post-training method called Canoe. The main idea behind Canoe is to synthesize easily verifiable short-form QA data and then leverage reinforcement learning (RL) with tailored rule-based rewards to improve the faithfulness of LLMs in both short-form and long-form generation tasks. Canoe firstly introduces Dual-GRPO, a variant of GRPO (Shao et al., 2024) that includes three carefully tailored rule-based RL rewards derived from synthesized short-form QA data, while optimizing both short-form and long-form response generation. For the provided contextual information and question, Dual-GRPO first prompts LLMs to produce a reasoning process, followed by a long-form answer composed of detailed and complete sentences, and finally a concise short-form answer in just a few words. In this way, we can assign different rewards to long-form and short-form responses, optimizing both simultaneously. Note that we assign accuracy rewards on generated short-form responses since the short-form QA task enables reliable rule-based verification of faithfulness. To overcome the problem of the faithfulness of the generated long-form responses being difficult to evaluate via rule-based verification (Zheng et al., 2025; OpenAI, 2025), we propose proxy rewards to evaluate it implicitly. Specifically, we construct the new input by replacing the given context with the generated long-form answer, then feed it to the LLMs to evaluate whether a long-form answer can drive the LLMs toward the correct short-form answer. If the generated long-form response enables LLMs to generate the correct final answer, this indicates that it remains context-faithful and contains easy-to-understand sentences that answer the question correctly. We also introduce format rewards to ensure more structured outputs and contribute to more stable training. To obtain the data used for training without human annotation, we collect head-relation-tail triples from the knowledge base, apply the advanced GPT-4o (OpenAI, 2023) to synthesize the question and contextual information, and use the tail entity from the triple as the answer to ensure the correctness. Moreover, we introduce four diverse QA tasks to ensure the complexity and diversity of the training data. Combined with the rule-based Dual-GRPO and data synthesis, Canoe can teach LLMs to remain context-faithful in both short-form and long-form generation tasks without relying on human annotations.
We evaluate the effectiveness of Canoe across 11 different downstream tasks, covering short-form and long-form generation tasks. Results show that Canoe significantly reduces faithfulness hallucinations. Specifically, Canoe significantly improves the overall score, e.g., 22.6% for Llama3-Instruct-8B. Meanwhile, Canoe surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score. To the best of our knowledge, these results are unprecedented for open-source models that do not rely on additional human annotations.
2 Related Work
Recently, the demand for utilizing LLMs to generate coherent text responses based on the provided contexts has continued to grow, particularly in text summarization and retrieval-augmented generation (RAG) scenarios. However, LLMs are often criticized for generating outputs that deviate from the provided contents, namely faithfulness hallucination (Li et al., 2022; Ji et al., 2023; Si et al., 2023; Huang et al., 2024). Many approaches have been proposed to improve the faithfulness of LLMs. The first line of work focuses on the inference stage of LLMs, such as designing prompts to encourage context integration (Zhou et al., 2023), improving context quality via explicit denoising (Xu et al., 2024a), and context-aware decoding to amplify contextual information (Shi et al., 2024). Although effective, these approaches primarily serve as a compensatory way rather than enabling the model to inherently learn to prevent generating unfaithful responses. Therefore, many studies attempt to apply post-training methods to improve the faithfulness. Bi et al. (2024) utilizes constructed faithful and unfaithful short-form completions and applies DPO to align LLMs to be context-faithful in short-form QA tasks. Huang et al. (2025) trains LLMs to discriminate between faithful and unfaithful responses in long-form QA tasks by unfaithful response synthesis and contrastive tuning. Duong et al. (2025) proposes a pipeline to generate a self-supervised task-specific dataset and applies preference training to enhance the faithfulness for a special task. However, these methods struggle to consistently improve the faithfulness of LLMs across various tasks, as these methods are designed for specific tasks. Thus, how to consistently improve the faithfulness of LLMs on different downstream tasks, including short-form and long-form generation tasks, still remains under-explored.
3 Methodology
Figure 2: An overview of Canoe framework. Canoe first synthesizes easily verifiable short-form QA data and then proposes the Dual-GRPO with designed rule-based rewards to improve the faithfulness of LLMs.
In this section, we will detail our proposed framework Canoe, which aims to teach LLMs to remain faithful across different tasks without human annotation. Specifically, we first synthesize easily verifiable short-form QA data and then propose the Dual-GRPO with designed rule-based rewards to improve the faithfulness of LLMs in both short-form and long-form response generation. We start with the introduction of the short-form data synthesis process, then a brief overview of RL protocol, and the tailored rule-based rewards used in the proposed Dual-GRPO training. An overview of the Canoe framework is presented in Figure 2.
3.1 Training Data Construction
Constructing high-quality and easily verifiable data is crucial for rule-based RL training (Shao et al., 2024). Inspired by knowledge base question generation (Cui et al., 2019; Guo et al., 2024), we attempt to collect triples from the knowledge base and use the advanced LLMs to synthesize the context and question. Concretely, we first collect about 30,000 head-relation-tail triples from Wikidata (Vrandečić and Krötzsch, 2014). Each collected triple $(h,r,t)$ includes a head entity $h$, a tail entity $t$, and the relation $r$ between two entities. Then we craft prompt templates and query the most advanced GPT-4o to synthesize the contextual information $c$ and question $q$ based on the triple $(h,r,t)$. We directly use the tail entity $t$ as the final answer $a$ to ensure the correctness and easy validation of the synthesized data. Each synthetic short-form QA sample $(c,q,a)$ consists of a contextual passage $c$, a question $q$, and a ground truth answer $a$. In this way, we can obtain short-form QA data that can be easily verified, thus we can utilize a rule-based RL method to optimize our LLMs to be more faithful. Meanwhile, to ensure the complexity and diversity of training data, we design four diverse QA tasks, including straightforward context, reasoning-required context, inconsistent context, and counterfactual context. The model is expected to answer the question by leveraging the information in the provided context.
Straightforward Context. A straightforward context means that the context clearly contains statements of the final answer. It requires models to accurately locate and utilize information from the context in order to answer questions. Specifically, we keep the original collected triple as input to query GPT-4o to synthesize the data $(c,q,a)$.
Reasoning-required Context. This context contains multiple related entities and relations, and requires models to answer multi-hop reasoning questions. Firstly, we construct a subgraph based on the sampled triples and extract 2, 3, 4-hop paths $[(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and employ the constructed paths to query GPT-4o to obtain the multi-hop context and question.
Inconsistent Context. This involves multiple randomly ordered contexts generated from different triples. This simulates noisy and inconsistent scenarios, where models need to detect inconsistencies and focus on useful and relevant contexts to answer the questions. We construct such a sample by combining the contexts from up to three QA samples.
Counterfactual Context. A counterfactual context contains statements that contradict common sense within the collected triples. Firstly, we replace the tail entity $t$ of the original collected triple with a similar but counterfactual entity $t^{cf}$. Then, we query GPT-4o to generate questions and counterfactual contexts to construct counterfactual samples. Unlike the aforementioned tasks, this task further highlights the importance of faithfulness for LLMs to answer the questions correctly, as it prevents models from depending on their learned factual knowledge to find the right answers.
By introducing four different tasks, we construct 10,000 QA pairs used for training without human annotation. These short-form QA data can be easily verified and include tasks varying in complexity, which can make rule-based RL training more efficient in improving the faithfulness of LLMs. More details can be found in the Appendix A, e.g., used prompts, data mixing recipes, and data statistics.
3.2 Reinforcement Learning Protocol
For RL training of LLMs, methods based on policy optimization, such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024), have been explored. Given the effectiveness of GRPO in training models and its advantages over PPO, e.g., eliminating the need for human-annotated preference data to train a reward model, we utilize GRPO to optimize and improve the faithfulness of the policy model $\pi_{\theta}$.
For each input, consisting of provided contextual information $c$, a natural language question $q$, the model generates a group of $G$ candidate answers, $\{o_{1},o_{2},\dots,o_{G}\}$. Each candidate is evaluated using a designed composite rule-based reward function to capture the end goal of faithfulness. GRPO leverages the relative performance of candidates within the group to compute an advantage $A_{i}$ for each output, guiding policy updates according to the following objective:
$\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{c,q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref})\right],\end{aligned}$
$\mathcal{L}_{i}&=\min\left(w_{i}A_{i},\text{clip}(w_{i},1-\epsilon,1+\epsilon)A_{i}\right),\end{aligned}$
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$, $\pi_{\theta_{\text{old}}}$ is the policy before the update, $\pi_{\text{ref}}$ is the reference policy (i.e., the initial model), $\epsilon$ and $\beta$ are hyperparameters controlling the update step and divergence regularization and $A_{i}$ is computed using the normalized reward within the group. We use synthesized short-form QA data as training data, which is easily verifiable, so that we can apply GRPO and train LLMs using the rule-based reward function. By generating multiple candidates per input, GRPO naturally accommodates the inherent challenges of utilizing the contextual information $c$ and answering the question $q$, e.g., LLMs may overly rely on the internal knowledge while disregarding provided contexts. Meanwhile, employing the rule-based GRPO removes the need for humans to annotate short-form and long-form preference data used for training the reward model.
3.3 Reward Design
Having a well-designed reward is key to the effectiveness of RL training (Du et al., 2025). To use easily verifiable short-form QA data to improve the faithfulness, the most intuitive reward would be the accuracy reward, which can check if the generated responses match the ground truth answers. However, in our early experiments, we found that relying solely on short-form QA data and accuracy rewards fails to enhance the faithfulness of long-form response generation, as the models may over-optimize short-form generation and learn a false pattern. For example, the tuned models tend to simply copy text spans from the context as answers and lose their ability to generate long-form responses. Unfortunately, directly evaluating the faithfulness of long, free-form responses via the rule-based verification continues to pose a significant and unresolved challenge.
Therefore, we propose Dual-GRPO, which includes a set of well-designed rewards that provide more harmonized guidance for optimizing LLMs to generate faithful responses. Unlike the original GRPO that over-optimizes short-form generation, we first prompt LLMs to generate both long-form and short-form responses, then assign different rewards to the two generated responses to improve the faithfulness of the two types of generation.
System Prompt and Rollouts. For the provided context and question, Dual-GRPO employs the designed system prompt that requires LLMs to produce a reasoning process, then a long-form answer composed of detailed and complete sentences, and finally a concise short-form answer in just a few words. For example, given the context, if the question is “What is the country of origin of Super Mario?”, the long answer could be “Super Mario originated from Japan.”, while the short answer could simply be “Japan”. In this way, we can assign different reward scores to long-form and short-form answers while optimizing them both at once. This system prompt also triggers zero-shot chain-of-thought reasoning in the policy model, which progressively improves as training advances to optimize for the reward. The system prompt used for Dual-GRPO rollouts is shown in the Appendix B.
Accuracy Reward for Short-form Response Generation. This reward directly assesses whether the generated short-form responses match the ground truth answers. We use the exact matching (EM) to measure accuracy, giving a score of 1 for a match and 0 for a mismatch. Thus, we can ensure that the generated short-form response correctly answers the question based on the context, making LLMs more faithful in short-form response generation.
Proxy Reward for Long-form Response Generation. Evaluating the faithfulness of the generated long-form responses via the rule-based verification remains challenging. This is because these long-form answers are often free-form, making rule-based verification ineffective (Zheng et al., 2025; OpenAI, 2025). Therefore, instead of directly evaluating the faithfulness of the long-form response, we propose a proxy reward to evaluate it implicitly, as the faithfulness of a long-form answer can be measured by its ability to drive the LLMs toward a correct short-form answer. Specifically, for each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and feed it to the LLM to check whether the LLM can produce the correct short-form answer based on $y_{lf}$. If the generated long-form response can enable the LLM to generate the correct answer, it indicates that the long-form response stays faithful to the context, contains complete and easy-to-understand sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct final answer, and a reward score of 0 for those that lead to an incorrect answer.
Format Reward. We also include a format reward that encourages adherence to a predefined output structure (e.g., using <think>, <long_answer>, and <short_answer> tags). Outputs that conform to this pattern receive a reward boost, thereby enhancing clarity and consistency. We use the string matching method to evaluate whether the generated responses adhere to the format, giving a score of 1 for a match and 0 for a mismatch.
Finally, we use the sum of these three rewards as the final composite reward. It enhances the efficacy of the rule-based RL training framework, guiding the model toward generating more faithful responses in both short-form and long-form tasks. More details are shown in the Appendix B.
4 Experiments
In this section, we conduct experiments and provide analyses to justify the effectiveness of Canoe.
Model Short-form Generation Tasks Long-form Generation Tasks Avg. Score
ConFiQA FiQA CNQ FaithEval FollowRAG XSum WikiLarge CLAPNQ
EM Acc EM Acc EM Acc Acc EM Acc FS FS FS Avg EM Avg Acc
The state-of-the-art LLMs
GPT-4o 31.5 42.7 66.8 79.6 43.4 55.9 47.5 42.2 57.8 80.7 88.1 70.3 58.8 65.3
GPT-4o mini 49.5 63.7 67.1 78.8 47.8 54.3 50.9 38.5 51.3 75.4 91.0 66.0 60.8 66.4
DeepSeek V3 49.5 58.6 67.0 76.5 54.6 67.3 51.0 37.7 55.2 82.8 85.6 71.0 62.4 68.5
Claude 3.7 Sonnet 26.0 36.0 56.4 72.2 41.4 65.0 45.6 36.3 53.7 78.3 81.7 68.3 54.3 62.6
[…]
DeepSeek R1 68.4 74.3 68.4 80.7 60.3 70.2 60.1 42.9 56.6 80.3 83.0 73.5 67.1 72.3
Claude 3.7 Sonnet-Thinking 27.1 38.7 59.5 76.7 42.1 67.0 57.0 38.8 55.3 79.0 81.4 72.2 57.1 65.9
LLaMA-3-Instruct Series
LLaMA-3-Instruct-8B 49.2 58.2 11.4 59.3 37.8 45.2 52.0 31.1 44.8 64.2 77.1 58.5 47.7 57.4
LLaMA-3-Instruct-70B 38.1 54.5 9.1 66.8 54.2 65.0 50.9 38.7 45.7 72.0 77.4 47.2 48.5 59.9
[…]
Context-DPO-8B 66.3 72.9 40.9 59.5 54.6 62.3 37.5 29.9 43.8 65.2 78.2 59.1 54.0 59.8
SCOPEsum-8B 35.7 64.6 7.1 68.7 33.8 60.6 55.7 30.1 46.2 70.3 80.3 59.8 46.6 63.3
Canoe-LLaMA-8B 73.5 80.9 82.7 84.9 66.7 73.4 74.6 40.9 51.7 74.4 84.4 64.9 70.3 73.6
$\Delta$ Compared to Vanilla. +24.3 +22.6 +71.3 +25.6 +28.9 +28.2 +22.6 +9.8 +6.9 +10.2 +7.3 +6.4 +22.6 +16.2
Qwen-2.5-Instruct Series
Qwen-2.5-Instruct-7B 52.5 61.0 13.2 68.4 55.3 68.2 56.1 32.6 45.3 63.4 57.8 61.2 49.0 60.2
Qwen-2.5-Instruct-14B 34.1 47.3 0.8 61.4 43.1 64.3 51.6 34.8 51.2 68.2 82.3 63.4 47.3 61.2
Qwen-2.5-Instruct-32B 44.5 66.4 39.2 81.1 37.7 66.4 47.0 33.9 53.1 20.2 57.7 31.7 39.0 52.9
Qwen-2.5-Instruct-72B 43.7 52.3 4.8 67.3 51.8 62.2 45.2 38.5 55.7 71.2 90.4 64.8 51.3 63.6
[…]
Context-DPO-7B 64.5 70.6 57.1 78.2 62.3 70.1 45.7 31.0 43.7 60.2 53.4 62.8 54.6 60.6
SCOPEsum-7B 39.3 47.9 12.9 60.9 50.2 55.3 52.3 30.6 46.0 68.3 72.0 63.2 48.6 58.2
Canoe-Qwen-7B 67.6 75.2 78.1 83.5 67.2 76.4 70.5 37.0 50.2 72.4 86.1 65.2 68.0 72.4
$\Delta$ Compared to Vanilla. +15.1 +14.2 +64.9 +15.0 +11.9 +8.2 +14.4 +4.4 +4.9 +9.0 +28.3 +4.0 +19.0 +12.3
[…]
$\Delta$ Compared to Vanilla. +51.6 +40.1 +87.0 +27.1 +38.7 +19.9 +15.8 +11.3 +3.4 +7.5 +8.8 +5.0 +28.2 +16.0
Table 1: Experimental results (%) on eleven datasets. The FollowRAG results represent the results averaged over these four open-domain QA datasets as shown in Table 7, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. Bold numbers indicate the best performance of models with the same model size. Avg EM/Acc represents the average score between short-form task metrics (EM/Acc) and long-form task metric FaithScore (FS).
4.1 Tasks and Datasets
To evaluate our method Canoe comprehensively, we select a range of downstream datasets, including short-form and long-form generation tasks.
Short-form Generation Tasks. For short-form generation tasks, we use two counterfactual QA datasets (ConFiQA (Bi et al., 2024) and CNQ (Longpre et al., 2021)), a multiple-choice questions dataset FaithEval (Ming et al., 2025), and a factual QA dataset FiQA (Bi et al., 2024) that is the factual version of ConFiQA. These datasets ensure the answers appear in the contexts to evaluate the faithfulness. We also evaluate our method on four open-domain QA datasets within the FollowRAG benchmark (Dong et al., 2024) to evaluate the abilities of LLMs in real-world RAG scenarios, including NaturalQA (Kwiatkowski et al., 2019b), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and WebQSP (Yih et al., 2016). In real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy. We evaluate models based on whether gold answers are included in the generated responses (i.e., Acc) following Asai et al. (2024) and exact matching (EM) for QA tasks. For multiple-choice questions, we follow Ming et al. (2025) and use keyword matching to verify the accuracy.
Long-form Generation Tasks. We include a text summarization task XSum (Narayan et al., 2018), a text simplification task WikiLarge (Zhang and Lapata, 2017), and a long-form QA task CLAPNQ (Rosenthal et al., 2025). To evaluate the faithfulness of generated long-form answers, called FaithScore (FS), we use MiniCheck (Tang et al., 2024) to check whether the model response is grounded in the provided context. MiniCheck is a state-of-the-art method to recognize if LLM output can be grounded in given contexts. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. We also query GPT-4o to evaluate the quality of generated responses, namely QualityScore.
More details are available in the Appendix C.
4.2 Baselines and Implementation Details
Baselines. We compare several baselines, including (1) Vanilla LLMs: including LLaMA-3-Instruct (Grattafiori et al., 2024) and Qwen-2.5-Instruct (Yang et al., 2024) of different sizes. We also conduct supervised fine-tuning on synthesized 10,000 short-form data as SFT baselines; (2) SOTA LLMs: We further evaluate the most advanced LLMs, including GPT-4o, GPT-4o-mini, OpenAI o1 (Jaech et al., 2024), Claude 3.7 Sonnet (Anthropic, 2025), Claude 3.7 Sonnet-Thinking, Deepseek R1, and Deepseek V3 (DeepSeek-AI et al., 2025a, b); (3) The Designed Methods to Improve Faithfulness of LLMs: Context-DPO (Bi et al., 2024) aligns LLMs through DPO with constructed faithful and unfaithful short-form answers, thus improving the faithfulness in short-form generation. SCOPE (Duong et al., 2025) introduces a pipeline to generate self-supervised task-specific data and applies preference training to enhance the faithfulness in a special task. We train it on the sampled training set of the summarization task XSum as SCOPEsum, regarding it as the method designed to improve the faithfulness of long-form response generation.
Implementation Details. Our main experiments are conducted on LLaMA-3-Instruct and Qwen-2.5-Instruct. More implementation details are shown in Appendix D, e.g., hyperparameters.
4.3 Main Results
Canoe Improves the Faithfulness of LLMs in Both Short-form and Long-form Response Generation. As shown in Table 1, Canoe shows consistent and significant improvements on 11 datasets measuring faithfulness. Canoe achieves substantial improvements in the overall score compared to original LLMs, e.g., 22.6% for Llama3-8B and 19.0% for Qwen2.5-7B in Avg EM score. Canoe also surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score (both Avg EM and Avg Acc scores). This shows that Canoe can effectively align LLMs to be context-faithful. Meanwhile, for real-world RAG scenarios, our proposed Canoe can also improve the performance even though the answer may not appear in the retrieved passages, and these passages are often noisy.
Model XSum WikiLarge CLAPNQ Avg
GPT-4o 98.5 97.5 81.2 92.4
[…]
LLaMA-3- Instruct-70B 86.2 83.0 30.1 66.4
Canoe-LLaMA-8B 85.8 87.8 65.5 79.7
[…]
Qwen-2.5-Instruct-14B 90.5 83.1 63.6 79.1
Qwen-2.5-Instruct-32B 90.3 83.9 58.6 77.6
Qwen-2.5-Instruct-72B 95.7 94.1 75.4 88.4
[…]
Canoe-Qwen-14B 91.9 89.7 73.5 85.0
Table 2: QualityScore on long-form generation tasks.
Model Acc EM
QA MR MC QA MR MC
GPT-4o 52.2 45.6 30.3 43.3 32.4 18.7
[…]
Canoe-LLaMA-8B 82.7 80.1 79.8 76.4 73.5 70.5
[…]
Qwen-2.5-Instruct-14B 62.4 44.9 34.7 44.7 34.3 23.3
Qwen-2.5-Instruct-32B 74.1 65.9 59.3 55.9 42.8 34.8
Qwen-2.5-Instruct-72B 63.3 50.3 43.3 54.3 42.2 34.7
[…]
Canoe-Qwen-14B 91.8 86.4 84.1 89.7 85.2 82.1
Table 3: Results (%) on three tasks in ConFiQA.
Figure 3: Model performance comparison on FaithEval in a closed-book QA setting and counterfactual context setting. Our models are colored in orange. We report the results from the chat version of LLaMA-3 and Qwen-2.5.
Canoe Maintains the Factuality of LLMs. We further evaluate whether Canoe will reduce the factuality of LLMs. Following Ming et al. (2025), we modify the original FaithEval and make it a closed-book QA setting, where no context is provided and LLMs need to give factual answers. In this case, the models rely entirely on their parametric knowledge of common facts, and we find that our proposed Canoe maintains the factuality compared to the untuned LLM as shown in Figure 3. However, when a new context with counterfactual evidence that contradicts the model’s parametric knowledge is introduced, performance declines sharply. For example, GPT-4o achieves 96.3% accuracy on factual closed-book QA task but only 47.5% on counterfactual QA task that evaluates the faithfulness of LLMs. This highlights that, unlike factuality, the faithfulness of LLMs is difficult to improve by simply scaling model parameters, which further indicates the necessity of a post-training method to improve faithfulness.
Canoe Improves the Quality of Long-form Response Generation. As shown in Table 2, we can find that our proposed Canoe also improves the quality of generations. This is because the proxy reward implicitly requires LLMs to generate easy-to-understand responses, which further optimizes the response quality. Canoe consistently improves the generation quality in the three long-form tasks, which illustrates the effectiveness of our method.
Canoe Enhances LLMs’ Reasoning in Short-form Response Generation. ConFiQA consists of three different tasks: question answering (QA), multi-hop reasoning (MR), and multi-conflicts reasoning (MC). QA focuses on the single-hop task with context containing one corresponding answer, while MR and MC involve multi-hop reasoning tasks with context containing one and multiple related counterfactual contexts, respectively. As shown in Table 3, Canoe not only improves the faithfulness in the single-hop QA task but also enhances the reasoning ability in reasoning tasks.
Canoe Mitigates Overconfidence Bias. For each model, we select a total of 110 unfaithful samples with the highest perplexity from the 11 datasets, 10 samples per dataset. Then we report the average perplexity score on these negative samples shown in Figure 4. We can find that Canoe produces the high perplexity scores, indicating low confidence scores, for these bad cases. This shows that Canoe mitigates overconfidence in these false statements.
Figure 4: The average perplexity score of 110 negative samples for each model from eleven datasets.
Model Short-form Tasks Long-form Tasks
EM Acc FaithScore QualityScore
Canoe-LLaMA-8B 67.7 73.1 74.6 79.7
-w/o. Dual-GRPO & Data Synthesis 36.3 51.9 66.6 64.3
[…]
-w/o. Inconsistent Context. 64.4 70.2 70.2 72.5
-w/o. Counterfactual Context. 62.6 67.8 69.7 73.7
Table 4: Results of ablation study. EM/Acc in short-form tasks represents the average score between QA metrics (EM/Acc) and the accuracy of FaithEval. N/A means a false generation pattern hacks this metric.
4.4 Analysis
Ablation Study. We conduct an ablation study in Table 4. The result reveals that our proposed Canoe (including Dual-GRPO and the designed short-form data synthesis) significantly improves the faithfulness of LLMs in both short-form and long-form generation. For Dual-GRPO, we observe that directly applying GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. We find that tuned models tend to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks (we show the case study in the Appendix F.5 to visually show this phenomenon). Thus, the generated responses do not contain syntactically and semantically complete sentences for long-form generation tasks, which leads to low QualityScore performance and also invalidates the metric used for evaluating faithfulness. We also explore the effectiveness of our training data construction strategy. For the designed QA tasks used to ensure the complexity and diversity of training data, we replace the designed Reasoning-required Context and Inconsistent Context samples with the same number of samples that contain Straightforward Context. We find that involving these more challenging instances can improve the effectiveness of RL training. We also replace the data points that contain Counterfactual Context with the same number of factual samples. The designed Counterfactual Context improves the final performance as it prevents models from depending on their learned factual knowledge to find the right answers.
Question: What do Greeks say when they break plates?
Qwen-2.5-Instruct-7B: Based on the information provided in the passages, there is no specific mention of what exactly Greeks say when they break plates during the plate-breaking ceremony. ()
[…]
Table 5: Case study from long-form QA task CLAPNQ. For different useful statements, we use different colors.
Case Study. We further conduct a case study in Table 4.4 to visually show the advantages of Canoe. Our method ensures the statements are faithful and comprehensive, and the text flows naturally.
Human Evaluation. Evaluating long-form generation tasks remains challenging (Li et al., 2024). Thus, we conduct human evaluation in the Appendix E to show the effectiveness of our method.
Discussion. We also discuss some possible concerns about Canoe in the Appendix F, e.g., the effect of the amount of synthesized data.
5 Conclusion
In this paper, we propose Canoe, a systematic post-training method for teaching LLMs to remain faithful in both short-form and long-form generation tasks without human annotations. By synthesizing diverse short-form QA data and introducing Dual-GRPO, a tailored RL method with three well-designed rule-based rewards, Canoe effectively improves the faithfulness of LLMs. We first synthesize short-form QA data with four diverse tasks to construct high-quality and easily verifiable training data without human annotation. We then propose Dual-GRPO, a rule-based RL method that includes three tailored rule-based rewards derived from synthesized short-form QA data, while optimizing both short-form and long-form response generation simultaneously. Experimental results show that Canoe consistently improves the faithfulness of LLMs across diverse downstream tasks.
Limitations
Although experiments have confirmed the effectiveness of the proposed Canoe, four major limitations remain. Firstly, Canoe synthesizes short-form QA data and uses the proposed Dual-GRPO to improve the faithfulness of LLMs in long-form response generation implicitly; thus, how to directly synthesize long-form data and improve the faithfulness remains under-explored. Meanwhile, the synthesized short-form QA data is single-turn; thus, exploring the synthesis of multi-turn QA data presents an attractive direction for future research. The motivation behind our work is to improve the faithfulness of LLMs without human annotation, but it is still worth exploring how to incorporate the existing manually labeled data to further improve the faithfulness of the model. Finally, while our method achieves strong results, exploring additional strategies, e.g., using cold-start to get a better initial policy model and improve the reward scores in training for better performance across different downstream tasks is also a promising direction.
Appendix
This appendix is organized as follows.
  • In Section A, we report the details of constructing training data, e.g., the used triples and introduction of four designed tasks.
  • In Section B, we go into detail about the proposed Dual-GRPO, including the system prompt and formal expressions of three different well-designed rewards.
  • In Section C, we show the details of evaluations, e.g., the introduction of the used benchmarks and evaluation prompts.
  • In Section D, we show the details of our implementation and training, e.g., hyperparameters and the used GPUs.
  • In Section E, we show the implementation details of human evaluation.
  • In Section F, we discuss some possible questions about the proposed Canoe. For example, we discuss the effect of the amount of synthesized short-form data for RL training.
Appendix A Training Data Details
A.1 Triples from Wikidata
To ensure the usability of the synthetic data and collected triples, we follow Bi et al. (2024) to collect entities corresponding to the top 1,000 most-visited Wikipedia pages from 2016 to 2023 and 41 relations selected by Bi et al. (2024) shown in Table 12. The most-visited Wikipedia pages are based on monthly page views and retain the most popular entities using criteria such as the number of hyperlinks. We finally collected 6,316 entities and 30,762 triples. We randomly select these triples to synthesize our training data, and finally construct 10,000 samples as the final training data.
A.2 Construction of Four Different Tasks
We design four different tasks to enhance the complexity and diversity of our training data. Meanwhile, we select GPT-4o-2024-08-06 to construct the contexts and questions.
Straightforward Context. As shown in Sec. 3.1, we keep the original collected factual triple as input to query GPT-4o to synthesize the data $(c,q,a)$. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 7 and Figure 8. We finally keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
Reasoning-required Context. We construct paths $[(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$ from a sub-graph; more details can be found in Sec. 3.1. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and use the constructed paths to query GPT-4o to obtain the multi-hop context and question. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 9 and Figure 10. We finally keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
Inconsistent Context. This involves multiple randomly ordered contexts generated from different triples. This simulates noisy and inconsistent scenarios, where models need to detect inconsistencies and focus on useful and relevant contexts to answer the questions. We construct such a sample by combining the contexts from up to three QA samples with reasoning-required context and use the original $t^{n}$ as the answer. In this way, we can obtain more complex samples than ones with the reasoning-required context. To avoid duplicating the 2,000 samples with the reasoning-required context collected above, we reconstruct the new samples with the reasoning-required context used to obtain the samples with the inconsistent context. We keep 1,000 such samples in the synthesized 10,000 training data, i.e., 10% of the data.
Counterfactual Context. A counterfactual context includes statements that go against common sense found in the collected triples. Specifically, we construct samples with counterfactual contexts below by modifying previously collected triples (of three types, including straightforward context, reasoning-required context, and inconsistent context). We replace the tail entity $t$ of the original collected triple with a similar but counterfactual entity $t^{cf}$, which is obtained by query GPT-4o using prompt “Generate me a noun for an entity that is similar to the {$t$} but different, and require the entity to exist in the real-world, please tell me the answer directly:”. Then, we query GPT-4o to generate questions and counterfactual contexts to construct counterfactual samples, using the counterfactual triples. The prompts used in constructing samples with counterfactual contexts are the same as the prompts used in constructing the three different tasks above. The reason we construct samples with counterfactual context in this way is that this prevents the model from learning the appropriate factual knowledge to answer the question correctly, rather than correctly exploiting the given contextual information. Therefore, we construct the same number of samples as the summed number of the three types above (including straightforward context, reasoning-required context, and inconsistent context), i.e., 5000 samples (50% of the data). Meanwhile, this task stresses the importance of keeping answers faithful in contexts, as it stops them from relying solely on the learned knowledge of LLMs to provide correct answers.
Type Num Avg Len
Straightforward Context. 2,000 186.3
Reasoning-required Context. 2,000 262.2
Inconsistent Context. 1,000 421.2
Counterfactual Context. 5,000 260.8
Table 6: Statistics of the training data. Num indicates the number of samples. Avg Len shows the average length of the samples, including the context and question.
A.3 Statistics
We show the statistics of the training data in Table 6. Even though the length of the data we synthesize is short, we find that our model can be generalized with consistently state-of-the-art results on a wide range of tasks with different input lengths by utilizing our proposed Dual-GRPO, e.g., long-form QA and RAG generation with long texts as inputs.
Appendix B Dual-GRPO Details
In this section, we give a more detailed introduction to our proposed Dual-GRPO, including the designed system prompt and formal expressions of three different rewards.
System Prompt. For the provided contextual information and question, Dual-GRPO employs the designed system prompt that requires LLMs to produce a reasoning process, then a long-form answer that consists of detailed and complete sentences, and finally a concise short-form answer in just a few words. In this way, we can assign different reward scores to long-form answers and short-form answers while optimizing them both at once. Meanwhile, this system prompt also triggers zero-shot chain-of-thought reasoning in the policy model, which progressively improves as training advances to optimize for the reward. We use the same system prompt to train both LLaMA and Qwen models. We show our used system prompt in Figure 11.
Accuracy Reward. For short-form generation, we directly assign the accuracy reward. Specifically, for the generated short-form response $y_{sf}$ based on the given context $c$ and question $q$, which is extracted from the whole generated response $y_{whole}$ via string matching, and the ground truth answer $y_{gt}$ from the synthesized training data, the accuracy reward $R_{\text{acc}}$ for the LLM $\theta$ can be calculated as: R_acc = {1 if ysf(c,q | θ) =ygt,0 otherwise.
We use the exact matching (EM) to measure accuracy, giving a score of 1 for a match and 0 for a mismatch. In this way, we can ensure that the generated short-form response correctly answers the question based on the given context, making LLMs more faithful in short-form response generation.
Proxy Reward. Instead of directly evaluating the faithfulness of the generated long-form response, we propose a proxy reward to evaluate it implicitly. Specifically, for each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and infer the LLM $\theta$ to determine whether the LLM can produce the correct short-form answer $y_{sf}$ based on $y_{lf}$ for the question $q$. Thus, the proxy reward $R_{\text{proxy}}$ can be calculated as: R_proxy = {1 if ysf(ylf,q| θ) =ygt,0 otherwise.
If the generated long-form response can help LLMs generate the correct answer, it indicates that the long-form response is faithful to the context, contains syntactically and semantically complete sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct answer, and a reward score of 0 for those that lead to incorrect answers.
Format Reward. To enforce the desired output format, we assign a reward on the whole generated response $y_{whole}$ to evaluate whether it contains the proper XML tags. We use three types of tags as shown in our system prompt, as shown in Figure 11, including <think>, <long_answer>, and <short_answer> tags. Formally, R_format = {1 if correct formatting is present,0 if incorrect formatting.
We use the string matching method to evaluate whether the responses adhere to the format.
Final Reward. Finally, we use the sum of these three rewards as the final composite reward $R_{\text{final}}$. This well-designed reward $R_{\text{final}}$ of Dual-GRPO enhances the efficacy of the rule-based RL training framework to guide the model toward generating more faithful responses in both short-form and long-form tasks. Formally, R_final = R_acc + R_proxy + R_format.
Finally, we use this reward $R_{\text{final}}$ to compute an advantage $A_{i}$ for each output, guiding policy updates according to the GRPO objective.
Potential Reward Hacking Concerns. In the early experiments, we have also tried adding the length reward for long-form responses (i.e., the content between <long_answer> and </long_answer> tags) to avoid the potential reward hacking, e.g., avoiding the policy model directly copying the given context as the long-form response, but found that the task performance does not have a significant difference.
Appendix C Evaluation Details
C.1 Datasets
ConFiQA (Counterfactual QA). This is a dataset that incorporates knowledge conflicts through counterfactual passages to evaluate the faithfulness of LLMs on short-form generation. ConFiQA consists of three tasks: QA (Question Answering), MR (Multi-hop Reasoning), and MC (Multi-Conflicts). QA features single-hop question-answering tasks with context containing one corresponding counterfactual, while MR and MC involve multi-hop reasoning tasks with context containing one and multiple related counterfactual contexts, respectively. ConFiQA contains 1,500 data points used for testing (500/500/500 from QA/MC/MR).
CNQ (Counterfactual QA). CNQ is constructed based on Natural Questions (Kwiatkowski et al., 2019a). In CNQ, the context is modified to support counterfactual answers following (Longpre et al., 2021). It contains 2,773 samples that incorporate counterfactual passages to evaluate the faithfulness of LLMs on short-form generation.
FaithEval (Counterfactual Multiple-choice QA). FaithEval is a novel and comprehensive benchmark tailored to evaluate the faithfulness of LLMs in contextual scenarios across three diverse tasks: unanswerable, inconsistent, and counterfactual contexts. We select the counterfactual task to evaluate the faithfulness of LLMs, which contains 1,000 multiple-choice QA samples curated based on ARC-Challenge (Clark et al., 2018).
FiQA (Factual QA). FiQA is a factual version of ConFiQA, which shares the same questions as ConFiQA but contains the factual contexts and answers. The contexts and answers are provided by Bi et al. (2024), thus we can evaluate the faithfulness of LLMs in factual short-form response generation. It contains 1,500 samples for evaluation.
FollowRAG (RAG Scenarios for short-form QA). FollowRAG aims to assess the model’s ability to follow user instructions in complex multi-document contexts. It consists of four well-known open-domain QA datasets for RAG scenarios, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. We utilize the provided passages in FollowRAG as context and original query (instead of the version with added instruction constraints proposed by Dong et al. (2024)) as questions. We also use the original answers to report the results. FollowRAG contains 2,800 samples used for testing (700/700/700/700 from NaturalQA/TriviaQA/HotpotQA/WebQSP). Different from short-form generation tasks that the contexts always contain answers, in real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy.
XSum (Summarization). Summarization is a content-grounded task where a model is provided a piece of text and tasked with synthesizing the most salient information within that text. XSum is a widely used dataset for text summarization, which consists of about 220,000 BBC articles as input documents. To facilitate our evaluation, we use the first 1,000 data points from the test set to evaluate our method.
WikiLarge (Simplification). Text simplification is a content-grounded task where a model is provided a piece of text and is tasked with paraphrasing it to make the text easier to read and understand. We use 1k instances sampled from the WikiLarge dataset as a test set, following Ravichander et al. (2025).
CLAPNQ (Long-form QA). CLAPNQ is a grounded long-form QA benchmark dataset for Retrieval Augmented Generation of LLMs. The answers are typically long, 2-3 sentences grounded on a single gold passage, in contrast to datasets based on machine reading comprehension, such as short-form Natural Questions, which are just a few words. CLAPNQ includes long answers with grounded gold passages from Natural Questions. We utilize the provided passages and questions from the dev set to evaluate the faithfulness of LLMs in long-form response generation for open-domain questions, which contains 600 data points.
C.2 Metrics and LLM-as-a-Judge
Metrics for Short-form Generation Tasks. We evaluate performance based on whether gold answers are included in the generated responses (i.e., Acc) following Asai et al. (2024) and exact matching (EM) for QA tasks. For multiple-choice questions in FaithEval, we use keyword matching to verify the accuracy, i.e., Acc.
Metrics for Long-form Generation Tasks. To evaluate the faithfulness of generated long-form answers, we use MiniCheck to check whether the model response is grounded in the provided context. MiniCheck is a state-of-the-art method to recognize if LLM output can be grounded in given contexts. We select the MiniCheck-FT5222https://huggingface.co/lytang/MiniCheck-Flan-T5-Large because it is the best fact-checking model, outperforming GPT-4o in evaluating the faithfulness. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. To evaluate the quality of the generated long-form responses for three different tasks (QualityScore), including summarization, simplification, and long-form QA, we design different prompts to query GPT-4o-2024-11-20 as a judge to get the quality scores. We report the average results of the quality score results by querying GPT-4o twice. The prompts for three tasks can be found in Figure 12, Figure 13, and Figure 14.
C.3 Baselines
For SOTA LLMs, we select the following versions of these models to report the results. Specifically, we use GPT-4o-2024-08-06 for GPT-4o, GPT-4o-mini-2024-07-18 for GPT-4o-mini, Claude 3.7 Sonnet-2025-02-19 for Claude 3.7 Sonnet and Claude 3.7 Sonnet-thinking, Deepseek R1 2025-01-20 for Deepseek R1, Deepseek V3 2024-12-26 for Deepseek V3, and o1-2024-12-17 for OpenAI o1. To get stable experimental results, we query these models twice and report the average results on each task. For the methods that are designed for improving the faithfulness, we reproduce their released code based on LLaMA-3-Instruct and Qwen-2.5-Instruct. For SCOPE, we train it on the 10,000 sampled training set of the summarization task XSum as SCOPEsum, which keeps the same number of data we used for training Canoe and provides a fair comparison.
C.4 Test-time Prompts
For baselines, the prompts for different tasks can be found in Figure 15, Figure 16, Figure 17, Figure 18, and Figure 19. To evaluate the factuality of LLMs, we modify the original FaithEval and make it a closed-book QA setting, and use the prompts shown in Figure 20. During the evaluation for Canoe, we apply the same system prompt during the Dual-GRPO training, and extract the content between <short_answer> and </short_answer> tags as the final answers for short-form generation tasks. Also, for long-form generation tasks, we extract the content between <long_answer> and </long_answer> tags as the final answers. We also find that the long-form responses generated by Canoe can provide correct answers in short-form generation tasks in the Appendix F.1. Thus, for real-world applications, we recommend using the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
Model HotpotQA NaturalQA TriviaQA WebQSP
EM Acc EM Acc EM Acc EM Acc
The state-of-the-art LLMs
GPT-4o 24.7 32.0 37.0 55.0 62.3 72.3 44.9 71.7
GPT-4o mini 18.0 26.2 35.0 48.2 59.5 65.5 41.4 65.3
DeepSeek V3 18.7 27.7 34.9 54.3 60.0 70.0 37.1 68.9
Claude 3.7 Sonnet 15.3 24.1 33.6 53.9 62.5 72.5 33.7 64.3
[…]
DeepSeek R1 26.0 29.3 38.7 52.9 68.0 73.0 38.9 71.3
Claude 3.7 Sonnet-Thinking 20.1 30.2 35.6 53.0 63.4 72.0 36.0 66.0
LLaMA-3-Instruct Series
LLaMA-3-Instruct-8B 13.0 18.2 31.0 40.3 45.5 60.2 35.0 60.4
LLaMA-3-Instruct-70B 24.1 28.7 36.5 45.3 63.0 66.6 31.3 42.1
[…]
Context-DPO-8B 10.1 16.7 23.4 37.8 53.3 62.3 32.8 58.3
SCOPEsum-8B 12.0 20.5 25.7 42.5 46.4 58.6 36.1 63.2
Canoe-LLaMA-8B 21.4 23.3 37.4 46.9 60.0 67.3 44.9 69.3
Qwen-2.5-Instruct Series
Qwen-2.5-Instruct-7B 14.0 17.6 32.2 42.3 50.3 62.3 33.9 58.8
Qwen-2.5-Instruct-14B 17.5 21.7 29.3 48.0 55.6 69.3 36.9 65.7
Qwen-2.5-Instruct-32B 16.5 24.6 26.3 50.2 50.0 70.7 42.7 66.7
Qwen-2.5-Instruct-72B 21.8 28.0 34.5 51.0 61.8 73.0 35.7 70.6
[…]
Context-DPO-7B 13.0 17.2 25.2 40.2 50.1 63.2 35.7 54.3
SCOPEsum-7B 12.5 19.5 27.2 43.5 48.4 60.1 34.2 60.7
Canoe-Qwen-7B 18.0 22.6 35.7 47.4 57.4 65.7 36.9 65.0
[…]
Table 7: Experimental results (%) on FollowRAG. Bold numbers indicate the best performance of models with the same model size.
C.5 More Detailed Experimental Results
FollowRAG contains four different QA datasets in RAG scenarios. We report the average results in Table 1. We show the more detailed results of FollowRAG in Table 7.
Figure 5: Human evaluation across four key dimensions.
Appendix D Implementations Details
We implement our method based on the RL framework open-r1 (Face, 2025). We use AdamW optimizer (Loshchilov and Hutter, 2019) to train our model, with a $1\times 10^{-6}$ learning rate, a batch size of 14 for 7B/8B models, and a batch size of 7 for the 14B model, steering the training across two epochs. We set the maximum input length for the models to 1,024 and the maximum generation length to 1,024. The number of generations $G$ during the RL training is set to 7, which is used in Eq. (1). We set 0.04 for $\beta$ used in Eq. (1). We set 0.2 for $\epsilon$ used for the clip shown in Eq. (2). We set 0.9 for temperature in RL training to generate responses. We conduct our experiments on NVIDIA A800-80G GPUs with DeepSpeed+ZeRO2 for 7B/8B models, DeepSpeed+ZeRO2+Offloading for the 14B model, and BF16. During the inference, we set 0.7 for temperature for the evaluation of our models and baselines. For each task, we infer the model twice and report the average scores as final results.
Appendix E Human Evaluation
We conduct a human evaluation on the 90 samples from long-form generation tasks, including 30/30/30 for summarization/simplification/long-form QA. We evaluate these samples across four key dimensions: readability, faithfulness, helpfulness, and naturalness. For each comparison, three options are given (Ours Wins, Tie, and Initial Model Wins), and the majority voting determines the final result. The participants follow the principles in Figure 21 to make the decision. We invite three Ph.D. students to compare the responses generated by the models. Before participants begin to make judgments, we describe the principles of our design in detail and ensure that each participant correctly understands the principles. If the final result can not be determined by majority voting, we will hold a discussion among the participants and vote on the result again. We compare two models, including Canoe-LLaMA-8B as our method and LLaMA-3-8B as the initial model. Shown in Figure 5, we can find that our method reduces faithfulness hallucinations and also ensures the response quality for three long-form generation tasks.
Model ConFiQA FiQA CNQ FaithEval HotpotQA NaturalQA TriviaQA WebQSP Avg
Acc Acc Acc Acc Acc Acc Acc Acc
The state-of-the-art LLMs
GPT-4o 42.7 79.6 55.9 47.5 32.0 55.0 72.3 71.7 57.1
GPT-4o mini 63.7 78.8 54.3 50.9 26.2 48.2 65.5 65.3 56.6
DeepSeek V3 58.6 76.5 67.3 51.0 27.7 54.3 70.0 68.9 59.3
Claude 3.7 Sonnet 36.0 72.2 65.0 45.6 24.1 53.9 72.5 64.3 54.2
[…]
DeepSeek R1 74.3 80.7 70.2 60.1 29.3 52.9 73.0 71.3 64.0
Claude 3.7 Sonnet-Thinking 38.7 76.7 67.0 57.0 30.2 53.0 72.0 66.0 57.6
LLaMA-3-Instruct Series
LLaMA-3-Instruct-8B 58.2 59.3 45.2 52.0 18.2 40.3 60.2 60.4 49.2
LLaMA-3-Instruct-70B 54.5 66.8 65.0 50.9 28.7 45.3 66.6 42.1 52.5
[…]
Context-DPO-8B 72.9 59.5 62.3 37.5 16.7 37.8 62.3 58.3 50.9
SCOPEsum-8B 64.6 68.7 60.6 55.7 20.5 42.5 58.6 63.2 54.3
Canoe-LLaMA-8B 80.9 84.9 73.4 74.6 23.3 46.9 67.3 69.3 65.1
   - Using Generated Long-form Responses. 92.3 95.5 81.6 78.2 32.7 59.3 74.1 79.1 74.1
   $\Delta$ Compared to Using Generated Short-from Response. +11.4 +10.6 +8.2 +3.6 +9.4 +12.4 +6.8 +9.8 +9.0
Qwen-2.5-Instruct Series
Qwen-2.5-Instruct-7B 61.0 68.4 68.2 56.1 17.6 42.3 62.3 58.8 54.3
Qwen-2.5-Instruct-14B 47.3 61.4 64.3 51.6 21.7 48.0 69.3 65.7 53.7
Qwen-2.5-Instruct-32B 66.4 81.1 66.4 47.0 24.6 50.2 70.7 66.7 59.1
Qwen-2.5-Instruct-72B 52.3 67.3 62.2 45.2 28.0 51.0 73.0 70.6 56.2
[…]
Context-DPO-7B 70.6 78.2 70.1 45.7 17.2 40.2 63.2 54.3 54.9
SCOPEsum-7B 47.9 60.9 55.3 52.3 19.5 43.5 60.1 60.7 50.0
Canoe-Qwen-7B 75.2 83.5 76.4 70.5 22.6 47.4 65.7 65.0 63.3
   - Using Generated Long-form Responses. 82.9 92.3 83.2 73.2 29.8 56.9 70.6 72.7 70.2
   $\Delta$ Compared to Using Generated Short-from Response. +7.7 +8.8 +6.8 +2.7 +7.2 +9.5 +4.9 +7.7 +6.9
[…]
   - Using Generated Long-form Responses. 89.8 94.4 87.1 70.6 30.0 58.0 73.1 76.6 72.5
   $\Delta$ Compared to Using Generated Short-from Response. +2.4 +5.9 +2.9 +3.2 +4.3 +6.4 +1.4 +7.3 +4.2
Table 8: Experimental accuracy score results (%) on short-form generation tasks. Bold numbers indicate the best performance among all the models.
Appendix F Discussion
F.1 Can Long-form Responses Generated by Canoe Provide Correct Answers in Short-form Generation Tasks?
This exploration is important because, in real-world applications, it is difficult to pre-determine whether to use generated short-form responses (i.e., the context between <short_answer> and </short_answer> tags) or long-form responses (i.e., the context between <long_answer> and </long_answer> tags) as answers to respond to user instructions. This contrasts with the evaluation of LLMs on different datasets, as described in the test-time strategies outlined in C.4. Therefore, we first explore whether the long-form responses generated by Canoe (i.e., the context between <long_answer> and </long_answer> tags) can provide correct answers in short-form generation tasks. As shown in Table 8, when evaluating the generated long-form responses that contain the free-form answers, the accuracy scores consistently increase in all the short-form generation tasks compared to using the generated short-form responses. It also indicates that the generated short-form responses maintain conciseness, which is important for measuring the EM score, but can slightly reduce the accuracy score. Therefore, in real-world applications, we can directly use the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only efficiently and faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
Model Accuracy Proxy Format
Canoe-LLaMA-8B 70.3 66.1 99.4
Canoe-Qwen-7B 64.1 63.4 99.9
Canoe-Qwen-14B 83.5 76.5 100.0
Table 9: Final rewards (%) in the RL training stage.
F.2 Final Rewards in the RL Training Stage
We show the final rewards in Table 9. We can find that models can easily learn the designed format, while accuracy and proxy rewards still remain challenging. Meanwhile, in the early stages of RL training, the format reward increases quickly and converges rapidly, and as training proceeds, the accuracy reward and the proxy reward gradually increase and eventually converge. This indicates that our well-designed training data construction strategy is effective and ensures the complexity and diversity, avoiding overfitting and reward hacking.
Model MultiFieldQA-zh DuReader VCSUM
LLaMA-3-Instruct-8B 80.1 65.2 42.2
Canoe-LLaMA-8B 88.2 75.3 65.2
[…]
Qwen-2.5-Instruct-14B 83.5 72.2 47.8
Qwen-2.5-Instruct-32B 85.1 77.2 52.7
Qwen-2.5-Instruct-72B 88.9 80.1 57.1
[…]
Canoe-Qwen-14B 93.2 84.3 70.4
Table 10: Results (%) on three Chinese datasets. Bold numbers indicate the best performance of models with the same model size.
F.3 Multilingual Transfer Ability and Context Length Generalization of Canoe
To further explore the multilingual transfer ability of Canoe, we further evaluate our model on the Chinese dataset. Specifically, we use the single-document QA dataset MultiFieldQA-zh (Bai et al., 2023), the multi-document QA dataset DuReader (He et al., 2018), and the summarization dataset VCSUM (Wu et al., 2023) within LongBench (Bai et al., 2023). Following Si et al. (2024) that utilizes the GPT-4 to evaluate the correctness of QA tasks and the faithfulness of the summarization task, we use the advanced LLM to evaluate these datasets. We use the same prompts for three tasks as Si et al. (2024) to query GPT-4o. The test-time prompts for these tasks can be found in Figure 23, Figure 24, and Figure 25. As shown in Table 10, we can find that our proposed Canoe also improves the faithfulness in Chinese datasets, indicating that our proposed method has a strong multilingual transfer ability. Meanwhile, these results also indicate that our method achieves better faithfulness even when our training data is short, i.e., Canoe guarantees a consistently strong performance of the LLM in long-context scenarios.
Figure 6: The Avg EM results (%) on 11 datasets with different numbers of synthesized short-form training data. We conduct the experiments based on LLaMA-3-Instruct-8B models.
F.4 Effect of the Amount of the Synthesized Short-form Data
To further explore the effect of the amount of the synthesized short-form data, we conduct the corresponding experiments as shown in Figure 6. We can find that increasing the number of training data improves the performance of our method, but when the number of data is greater than 10,000, the performance will be stable. For the models with different sizes, the optimal amount of training data may be different, e.g., the larger models may need more training data to achieve optimal performance. Fortunately, due to our training data construction strategy, we can simply scale and synthesize training data without human annotation.
Long-form QA
Inputs:{}
Question: What is the type and definition of this dance?
[…]
Simplification
Inputs: The Pokémon series began with the release of Pokémon Red and Green for the Game Boy in Japan. When these games proved extremely popular, an enhanced blue version was released sometime after, and the blue version was reprogrammed as Pokémon Red and Blue for international release.
[…]
Summarization
Inputs: {}
[…]
Table 11: Case study between the models trained via GRPO and our model for long-form generation tasks. Models are implemented on LLaMA-3-Instruct-8B.
F.5 Case Study between GRPO and the proposed Dual-GRPO
We find that directly applying GRPO instead of our proposed Dual-GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. The used system prompt for applying GRPO can be found in Figure 22. Shown in Table F.4, we can find that the tuned model GRPO-LLaMA-8B tends to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks. However, when we apply Dual-GPRO to our synthesized data, we find that trained models can generate fluent and complete sentences. Thus, Dual-GRPO not only improves the faithfulness of LLMs in two types of response generation but also ensures the utility of models.
Relation Description
P6 head of government
P17 country
P26 spouse
P27 country of citizenship
P30 continent
P35 head of state
P36 capital
P37 official language
P38 currency
P39 position held
P50 author
P54 member of sports team
P57 director
P86 composer
P101 field of work
P103 native language
P108 employer
P112 founder
P127 owned by
P136 genre
P1376 capital of
P140 religion
P155 follows
P159 headquarters location
P166 award received
P170 creator
P172 ethnic group
P175 performer
P178 developer
P264 record label
P276 location
P286 head coach
P407 language of work or name
P413 position played
P463 member of
P488 chairperson
P495 country of origin
P641 sport
P800 notable work
P937 work location
P169 chief executive officer
Table 12: Manually selected relations that are used to construct training data. We utilize the same manually selected relations as Bi et al. (2024).
Prompt for question generation for the samples with straightforward context. [Instructions] You are a sophisticated question generator. Given a triple {$(h,r,t)$} collected from Wikidata, generate a question that asks about the final tail entity {$t$} using the head entity {$h$} and the relation {$r$}. Directly give me the generated question:
Figure 7: Prompt for question generation for the samples with straightforward context.
Prompt for context generation for the samples with straightforward context. [Instructions] You are a sophisticated context generator. Given a triple {$(h,r,t)$} collected from Wikidata, generate a brief description of the head entity {$h$}, approximately 150 words long. Ensure the tail entity {$t$} and relation {$r$} are accurately mentioned in the generated description. Directly give me the generated context:
Figure 8: Prompt for context generation for the samples with straightforward context.
Prompt for question generation for the samples with reasoning-required context. [Instructions] You are a sophisticated question generator. Given a chain of triples {[…]} collected from Wikidata, generate a question that asks about the final tail entity {$t$} using the head entity {$h$} and the relation {$r$}. Do not include any bridge entities in the question; instead, phrase the question as if directly asking about the relationship from the head entity to the tail entity Directly give me the generated question:
Figure 9: Prompt for question generation for the samples with reasoning-required context.
Prompt for context generation for the samples with reasoning-required context. [Instructions] You are a sophisticated context generator. Given a chain of triples {[…]} collected from Wikidata, generate a brief description of the head entity {$h$}, approximately {150*n} words long. Ensure the tail entity {$t$} and relation {$r$} are accurately mentioned in the generated description. Directly give me the generated context:
Figure 10: Prompt for context generation for the samples with reasoning-required context.
System prompt for Dual-GRPO. A conversation between User and Assistant. The user gives an instruction that consists of two parts: a passage and the actual instruction, separated by two newline characters. The passage is provided within <context> and </context> tags. The Assistant needs to refer to the given passage and complete the instruction. The Assistant solves the question by first thinking about the reasoning process internally, according to the given passage, and then providing the response. The response must be structured and include the following three sections, clearly marked by the respective tags: - Reasoning Process: Explain your thought process or logical steps to derive the answer. Enclose this within <think> and </think> tags. - Long Answer: Provide a long response that consists of syntactically and semantically complete sentences to answer the question. Enclose this within <long_answer> and </long_answer> tags. - Short Answer: Present a concise response that directly answers the question. Enclose this within <short_answer> and </short_answer> tags. Format your response exactly as follows: <think> reasoning process here. </think> <long_answer> detailed answer here. </long_answer> <short_answer> the concise answer here. </short_answer>.
Figure 11: System prompt for Dual-GRPO.
Prompt used to calculate quality score for text summarization. You are asked to evaluate the quality of the AI assistant’s generated summary as an impartial judge, and your evaluation should take into account factors including readability (whether the summary is clear and easy to understand) and coherence (whether the assistant’s summary is logical and orderly). Read the AI assistant’s summary and input passages, and give an overall integer rating in on a scale of 1 to 5, where 1 is the lowest and 5 is the highest based on the evaluation criteria, strictly in the following format:“[[rating]]”, e.g. “[[5]]”. Input Passages: {} Assistant’s summary:{} Rating:
Figure 12: Prompt used to calculate quality score for text summarization.
Prompt used to calculate quality score for text simplification. You are asked to evaluate the quality of the AI assistant’s generated text simplification as an impartial judge, and your evaluation should take into account factors including readability (whether the simplification is clear and easy to understand) and coherence (whether the assistant’s simplification is logical and orderly). Read the AI assistant’s simplified version and the original text, and give an overall integer rating on a scale of 1 to 5, where 1 is the lowest and 5 is the highest based on the evaluation criteria, strictly in the following format: “[[rating]]”, e.g. “[[5]]”. Original text: {} AI assistant’s simplification: {} Rating:
Figure 13: Prompt used to calculate quality score for text simplification.
Prompt used to calculate quality score for long-form QA. You are asked to evaluate the quality of the AI assistant’s generated long-form answer as an impartial judge, and your evaluation should take into account factors including readability (whether the answer is clear and easy to understand) and coherence (whether the answer is logical and well-organized). Read the AI assistant’s long-form answer and the original question, and give an overall integer rating on a scale of 1 to 5, where 1 is the lowest and 5 is the highest, based on the evaluation criteria, strictly in the following format: “[[rating]]”, e.g., “[[5]]”. Question: {} Assistant’s long-form answer: {} Rating:
Figure 14: Prompt used to calculate quality score for long-form QA.
Test-time prompt used for short-form QA tasks. Passages: {} Refer to the passages above and answer the following question with just a few words. Question: {} Answer:
Figure 15: Test-time prompt used for short-form QA tasks.
Test-time prompt used for multiple-choice QA task. Passages: {} Refer to the passages above and answer the following question with just a few words. Question: {} Please select the correct option according to the question, and output the option letter (e.g. A/B/C/D): Options: {} Answer:
Figure 16: Test-time prompt used for multiple-choice QA task.
Test-time prompt used for text summarization. Passage: {} Refer to the passage above and provide a summary as the response. Summary:
Figure 17: Test-time prompt used for text summarization.
Test-time prompt used for text simplification. Passage: {} Refer to the passage above and simplify it to improve its readability, ensuring its core meaning remains intact. Please provide only the simplified text as the response. Simplified text:
Figure 18: Test-time prompt used for text simplification.
Test-time prompt used for long-form QA task. Passage: {} Refer to the passages above and answer the following question. Question: { }
Figure 19: Test-time prompt used for long-form QA task.
Test-time prompt used for FaithEval in closed-book QA settings. Question: {} Please select the correct option according to the question, and output the option letter (e.g. A/B/C/D): Options: {} Answer:
Figure 20: Test-time prompt used for FaithEval in closed-book QA settings.
The principles of human evaluation for long-form responses generation. You are asked to evaluate the responses generated by different models. You should choose the preferred responses according to the following perspectives independently:
1. Readability: Whether the response is clear and easy to understand? 2. Faithfulness: Whether the response is faithful to the context and the information can be grounded in the provided context. 3. Helpfulness: Whether the response provides useful information and follows the instructions from users? 4. Naturalness: Whether the response sounds natural and fluent? Finally, please make a decision among the 3 opinions, including Win, Tie, and Loss.
Figure 21: The principles of human evaluation for long-form responses generation.
System prompt for GRPO in the ablation study. A conversation between User and Assistant. The user gives an instruction that consists of two parts: a passage and the actual instruction, separated by two newline characters. The passage is provided within <context> and </context> tags. The Assistant needs to refer to the given passage and complete the instruction. The Assistant solves the question by first thinking about the reasoning process internally, according to the given passage, and then providing the response. The response must be structured and include the following two sections, clearly marked by the respective tags: - Reasoning Process: Explain your thought process or logical steps to derive the answer. Enclose this within <think> and </think> tags. - Answer: Present a concise response that directly answers the question. Enclose this within <answer> and </answer> tags. Format your response exactly as follows: <think> reasoning process here. </think> <answer> answer here. </answer>.
Figure 22: System prompt for GRPO in the ablation study.
Test-time prompt used for MultiField-zh. {CJK} UTF8gbsn 阅读以下文字并用中文简短回答:{} 现在请基于上面的文章回答下面的问题,只告诉我答案,不要输出任何其他字词。 问题:{} 回答:
Figure 23: Test-time prompt used for MultiField-zh.
Test-time prompt used for DuReader. {CJK} UTF8gbsn 请基于给定的文章回答下述问题。 文章:{} 问题:{} 回答:
Figure 24: Test-time prompt used for DuReader.
Test-time prompt used for VCSUM. {CJK} UTF8gbsn 下面有一段会议记录,请你阅读后,写一段总结,总结会议的内容。 会议记录:{} 会议总结:
Figure 25: Test-time prompt used for VCSUM.